Show all the solutions. Round off answers to two decimal places. Be consistent with the units.

4. A 40-ohm resistor is to be wound from 0.0001 cm2 platinum wire. How many meters of wire are needed? (Refer to the table of resistivities under the theory for the value of resistivity of  platinum)


5. A silver wire 2 m long is to have a resistance of 0.5 ohm. What should be its

 cross-sectional area? (Refer to the table of resistivities under the theory for the value of resistivity of silver)


1
Expert's answer
2021-04-23T07:34:39-0400

Answer

4.

Resistance is given by

"R=\rho \frac{l}{A}"


"40=(1.06\times 10^{-7})\frac{l}{10^{-8}}\\l=3.77m"



5.

"R=\rho \frac{l}{A}"


So

"0.5=(1.59\times10^{-9})\frac{(2)}{A}\\A=6.36\times10^{-9}m^2"